Simplify:
(x+2)(x-1)

Answer by algebrahouse.com(1659)

"Simplify:
(x+2)(x-1)"

I assume you mean use the foil method.

(x + 2)(x - 1)
= x(x) + x(-1) + 2(x) + 2(-1) {used foil method}
= x^2 - x + 2x - 2 {multiplied through}
= x^2 + x - 2 {combined like terms}
www.algebrahouse.com